X=2/5 y+1 how do you get this in y intercept form?
Answer by Cromlix(4381)

x = 2/5y + 1
-2/5y = -x + 1
2/5y = x - 1
multiply thro' by 5/2
y = 5/2x - 5/2
To remove the fractions
multiply thro' by 2
2y = 5x - 5
Hope this helps.
:-)